OpenStudy (anonymous):

the second term in a geometric sequence is 12 . the fourth term in the same sequence is 4/3. what is the common ratio in this sequence

OpenStudy (anonymous):

t2=12 t4=4/3 4/3 divided by 12=1/9. then, t3/t1 (because in every equation, its n-1) is t2. so t2=1/9. what's the square root of 1/9? 1/3. so that's your common ratio. (sorry if my conclusion didn't make much sense)